Using Substitution, solve.
x+y=4
y=2x+1

Answer by pie90010(44)

So it depends which variable you want to use, ill use x in this case
take the equation x+y=4 and isolate the x so its
x = (4 - y)
then in the other equation, replace x with (4-y)
y = 2(4 - y) + 1
Now you have only one variable :D
y = 8 - 2y + 1
combine like terms, yada yada
3y = 9
y = 3

Now just plug the number in to the original equation to get X
x + 3 = 4
X = 1

the answer would be (1,3)